Sol Summary – December 13th

Summary Title:  In the final stretch (and clouds)

Mission Status:

We are getting close to the end of our mission, and the focus of the crew is maximizing the time we have left to complete our science objectives.

Today in the Matryoshka EVA series. Crewmembers Trisha Randazzo, Joshua Hunt, and Willie Schumann traveled to “Matrysoshka Site 6” which was located in the far north at the begging of Lith Canyon. After a long drive to the site with

We attempted to do solar observation tonight but were unfortunately stymied by extensive cloud cover all day long today.

This evening our team will focus on Anesthesia Experiment runs.  Once complete our team is currently hoping the clouds clear to allow us to watch the Geminid meteor shower which peaks tonight.  Shannon also allowed us to assemble her old telescope for use tonight.  Fingers crossed the clouds clear!

We preparing for our last ‘science day’ on Thursday, as we finish up both Matryoshka, Bricks, and Anesthesia experiments.  Friday is being held in reserve and is also allocated to cleaning the HAB and writing up lessons learned for the next crew.  The end is fast approaching!

Thanks for your help!

Sol Activity Summary:

  1. Matryoshka EVA Site #6
  2. ‘Bricks’ communication experiment
  3. Exercise
  4. General HAB Cleaning / Maintenance / Reports
  5. Telescope Assembly
  6. Anesthesia Experiment Runs this evening
  7. (Hopefully)  Geminid Observation!

Look Ahead Plan:

  1. Matryoshka EVA Site #7a
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. Anesthesia Experiment Runs
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports

Anomalies in work:

1.    HAB Leak from under EVA door.  No crew has showered yet today and a significant leak was observed by returning EVA crew.  It is likely this leak is coming from the Kitchen as the dishes were washed immediately prior to the crew returning.

2.  Power System Misconfiguration ongoing.

Weather: Mild and Cloudy

Crew Physical Status: All crew in good physical health.

EVA: See Above

Reports to be filed:



EVA Request

Sol Summary



Support Requested:

1.  We have an extra CO detector that is marked ‘For Greenhab’.  The current detector in Greenhab is both a Smoke Alarm / CO Sensor.  Please advise whether you would like us to install a redundant CO Sensor in the Greenhab or return it to the HAB and mark it as a replacement part.

2.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken
    16. 1 Can Strawberries
    17. 1 Box Granola
    18. 1 Bottle Soy Sauce
Copyright © The Mars Society. All rights reserved. | Main Site